# Quantum Mpemba Effect in Dissipative Spin Chains at Criticality

### v1

@@ I Introduction
The QME has attracted increasing attention following its realization in engineered quantum platforms [14, 15, 16]. These observations have led to intensive theoretical efforts to understand the underlying mechanisms in various physical regimes, including quantum integrable systems [17, 18, 19, 20, 21], disordered systems [22, 23, 24], random quantum circuits [25, 26], and other settings  [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. A natural theoretical framework for describing QME and its inverse is through open quantum systems. The experimental realization of open quantum systems in a variety of controllable platforms [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50] has reinvigorated research into dissipative quantum dynamics governed by master equations. They provide means to characterize QME by analyzing relaxation pathways from different initial states and investigating how decoherence shapes the approach to steady state. In open quantum systems governed by Lindblad dynamics, the QME manifests as a non-monotonic dependence of relaxation times on initial conditions. This behavior is rooted in the spectral structure of the Liouvillian superoperator. Depending on how the initial state overlaps with slow-decaying modes, relaxation can be anomalously accelerated.
@@ III Strong quantum Mpemba effect in at critical points
Figure 4: Dynamical evolution near the antiferromagnetic quantum critical point of the XXZ model. The initial states are chosen as the highest excited state (blue line) and a finite negative-temperature state (red line), with system size $L=8$. (a) For $\Delta=-1.1$, no QME is observed. (b) At the critical point $\Delta=-1$, sQME emerges. (c) For $\Delta=-0.9$, no QME is observed. The blue line denotes the evolution from initial state with temperature $T=0^{-}$ and the red one is T=5T=-5.
@@ V Conclusion and outlook
In this work, we have investigated the QME in open quantum spin chains subject to dephasing dissipation. By tracking the complete heating dynamics from thermal states at different initial temperatures, we identified the emergence of a sQME in the vicinity of quantum critical points. In this strong form, a state prepared farther from equilibrium relaxes faster than any state initialized closer to the steady state-a marked departure from the wQME, which only requires a non-monotonic relaxation profile. By analyzing the spectral properties of the Liouvillian superoperator and the overlap distributions of different initial states with its eigenmodes, we established the physical mechanism underlying the emergence of sQME. In the XXZ chain, we demonstrated that the sQME arises precisely at the ferromagnetic ($\Delta=1$) and antiferromagnetic ($\Delta=-1$) quantum critical points, where the zero-temperature state relaxes faster than all finite-temperature states despite being farther from equilibrium. We further extended our analysis to the frustrated $J_{1}$$J_{2}$ XXZ model, identifying a robust sQME in the parameter regime $J_{2}/J_{1}>-0.25$ at $\Delta=1$, and revealed its remarkable sensitivity to the frustration-induced phase transition at $J_{2}/J_{1}=-0.25$. Across all cases, we found that the sQME survives exclusively along the isotropic line $\Delta=1$, while deviations in anisotropy, system size, or temperature reduce it to a wQME. Our findings highlight the intimate connection between non-equilibrium relaxation dynamics and underlying quantum criticality. They also suggest that the QME provides a novel dynamical probe for detecting critical points and characterizing frustrated quantum phases. On the experimental side, the models we studied can be realized in ultracold atomic gases, trapped ions, and superconducting qubits, and dephasing-type dissipation has already been engineered in such platforms, making the observation of the QME experimentally feasible.
@@ Acknowledgements
The work is supported by the National Natural Science Foundation of China (Grant No. 12304290) and the Fundamental Research Funds for the Central Universities.

### v2

@@ I Introduction
The QME has attracted increasing attention following its realization in engineered quantum platforms [14, 15, 16]. These observations have led to intensive theoretical efforts to understand the underlying mechanisms in various physical regimes, including quantum integrable systems [17, 18, 19, 20, 21], disordered systems [22, 23, 24], random quantum circuits [25, 26, 27], and other settings  [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. A natural theoretical framework for describing QME and its inverse is through open quantum systems. The experimental realization of open quantum systems in a variety of controllable platforms [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52] has reinvigorated research into dissipative quantum dynamics governed by master equations. They provide means to characterize QME by analyzing relaxation pathways from different initial states and investigating how decoherence shapes the approach to steady state. In open quantum systems governed by Lindblad dynamics, the QME manifests as a non-monotonic dependence of relaxation times on initial conditions. This behavior is rooted in the spectral structure of the Liouvillian superoperator. Depending on how the initial state overlaps with slow-decaying modes, relaxation can be anomalously accelerated.
@@ III Strong quantum Mpemba effect in at critical points
Figure 4: Dynamical evolution near the antiferromagnetic quantum critical point of the XXZ model. The initial states are chosen as the highest excited state (blue line) and a finite negative-temperature state (red line), with system size $L=8$. (a) At the critical point $\Delta=-1$, sQME emerges. (b) For $\Delta=-1.1$, no QME is observed. (c) For $\Delta=-0.9$, no QME is observed. The blue line denotes the evolution from initial state with temperature $T=0^{-}$ and the red one is T=5T=-5.
@@ V Conclusion and outlook
In this work, we have investigated the QME in open quantum spin chains subject to dephasing dissipation. By tracking the complete heating dynamics from thermal states at different initial temperatures, we identified the emergence of a sQME in the vicinity of quantum critical points. In this strong form, a state prepared farther from equilibrium relaxes faster than any state initialized closer to the steady state–a marked departure from the wQME, which only requires a non-monotonic relaxation profile. By analyzing the spectral properties of the Liouvillian superoperator and the overlap distributions of different initial states with its eigenmodes, we established the physical mechanism underlying the emergence of sQME. In the XXZ chain, we demonstrated that the sQME arises precisely at the ferromagnetic ($\Delta=1$) and antiferromagnetic ($\Delta=-1$) quantum critical points, where the zero-temperature state relaxes faster than all finite-temperature states despite being farther from equilibrium. We further extended our analysis to the frustrated $J_{1}$$J_{2}$ XXZ model, identifying a robust sQME in the parameter regime $J_{2}/J_{1}>-0.25$ at $\Delta=1$, and revealed its remarkable sensitivity to the frustration-induced phase transition at $J_{2}/J_{1}=-0.25$. Across all cases, we found that the sQME survives exclusively along the isotropic line $\Delta=1$, while deviations in anisotropy, system size, or temperature reduce it to a wQME. Our findings highlight the intimate connection between non-equilibrium relaxation dynamics and underlying quantum criticality. They also suggest that the QME provides a novel dynamical probe for detecting critical points and characterizing frustrated quantum phases. On the experimental side, the models we studied can be realized in ultracold atomic gases, trapped ions, and superconducting qubits, and dephasing-type dissipation has already been engineered in such platforms, making the observation of the QME experimentally feasible.
@@ Acknowledgements
The work is supported by the National Key Re- search and Development Program of China (Grants No. 2023YFA1406704 and No. 2022YFA1405800) and National Natural Science Foundation of China (Grant No. 12304290 and No. 12474496). LP also acknowledges support from the Fundamental Research Funds for the Central Universities.